OpenStudy (anonymous):

How to solve sqrt(x -1) = 0??

OpenStudy (anonymous):

I can't believe I forgot how to. I know the answer is one but how do you work it?

OpenStudy (xapproachesinfinity):

where did you stuck

OpenStudy (anonymous):

@xapproachesinfinity the answer is one but can you tell me how to solve it? Do you square it or just add 1 to both sides??

OpenStudy (xapproachesinfinity):

you need to square root both sides\( \sqrt{x-1}=0 ====>(\sqrt{x-1})^2=0^2==>x-1=0\)

OpenStudy (anonymous):

look first you square both sides to get (x-1 = 0) then add 1 to both sides to get (x=1

OpenStudy (anonymous):

is it cleear ???

OpenStudy (xapproachesinfinity):

How about this one \(\sqrt{(x+1)^2}=?\)

OpenStudy (anonymous):

Okay. That's what I was asking, I just wanted reassurance. No need for hasty or rude comments. Thanks @xapproachesinfinity.

OpenStudy (xapproachesinfinity):

yw! there was no rude comment! you are talking about the other guy? just don't pay attention to him!

OpenStudy (anonymous):

what ???? i didnt make any rude comments !!!

OpenStudy (xapproachesinfinity):

Not you @Hamoody1996! there was a guy who made some comment didn't even bother to read his comment, but it was of course out of topic and rude i guess

OpenStudy (xapproachesinfinity):

So @xuvaze what do you think about the root i put!

OpenStudy (anonymous):

ohhh i didnt even noticed his comment

OpenStudy (xapproachesinfinity):

it think it was deleted! Ate any rate, we have no concern about that!